OpenStudy (anonymous):

A trunk w/ a mass of 100kg is pulled 10m across a floor by a rope making an angle of 30 degree w/ the horizantal.if the coeffient of sliding frictions is 0.2.find a. the force of the rope b.the work done

OpenStudy (anonymous):

Fcos30-.2*(100-Fsin30)>=0 F=20.70

OpenStudy (anonymous):

w=20.70*cos30*10=179.29